Thompson Industrial Products Inc. (TIPI) is a diversified industrial-cleaner processing company. The company’s Dargan plant produces two products: a table cleaner and a floor cleaner from a common set of chemical inputs (CDG). Each week, 904,500 ounces of chemical input are processed at a cost of $210,300 into 603,000 ounces of floor cleaner and 301,500 ounces of table cleaner. The floor cleaner has no market value until it is converted into a polish with the trade name FloorShine. The additional processing costs for this conversion amount to $255,500.

FloorShine sells at $22 per 30-ounce bottle. The table cleaner can be sold for $17 per 25-ounce bottle. However, the table cleaner can be converted into two other products by adding 301,500 ounces of another compound (TCP) to the 301,500 ounces of table cleaner. This joint process will yield 301,500 ounces each of table stain remover (TSR) and table polish (TP). The additional processing costs for this process amount to $100,000. Both table products can be sold for $14 per 25-ounce bottle.

The company decided not to process the table cleaner into TSR and TP based on the following analysis.

Process Further
Table
Cleaner
Table Stain
Remover (TSR)
Table
Polish (TP)
Total
Production in ounces 301,500 301,500 301,500
Revenues $205,020 $168,840 $168,840 $337,680
Costs:
   CDG costs 70,100 * 52,575 52,575 105,150 **
   TCP costs 0 50,000 50,000 100,000
     Total costs 70,100 102,575 102,575 205,150
Weekly gross profit $134,920 $66,265 $66,265 $132,530


*If table cleaner is not processed further, it is allocated 1/3 of the $210,300 of CDG cost, which is equal to 1/3 of the total physical output.
**If table cleaner is processed further, total physical output is 1,206,000 ounces. TSR and TP combined account for 50% of the total physical output and are each allocated 25% of the CDG cost.

Determine if management made the correct decision to not process the table cleaner further by doing the following.

(1) Calculate the company’s total weekly gross profit assuming the table cleaner is not processed further.

Total weekly gross profit $

Entry field with incorrect answer


(2) Calculate the company’s total weekly gross profit assuming the table cleaner is processed further.

Total weekly gross profit $

Entry field with incorrect answer


(3) Compare the resulting net incomes and comment on management’s decision.

Management made the

Entry field with correct answer rightwrong

decision by choosing to not process table cleaner further.


Using incremental analysis, determine if the table cleaner should be processed further. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Don’t Process
Table Cleaner
Further
Process
Table Cleaner
Further
Net Income
Increase
(Decrease)
Incremental revenue $

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer

Incremental costs

Entry field with incorrect answer

Entry field with incorrect answer

Entry field with incorrect answer

   Totals $

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer
Answer #1

Allocated cost is irrelevant since already incurred

1.Weekly gross profit = 205020 - 0 = $205,020

2.If processed further = 337,680-100,000 = $237,680

Made wrong decision

Don’t Process

Process

Net Income Increase

Incremental Revenue

          205,020

337,680

132,660

Incremental costs

   0   

100,000

(100,000)

Totals

205,020

237,680

32,660